# Optimal Switching in Networked Control Systems: Finite Horizon

### 29 Mar 2026 v1

@@ 4.2 The Symmetry Assumption in Prior Work
In this section, we begin by reviewing common triggering mechanisms in the literature that motivate the symmetric policy assumption in the delay-free setting. As noted in Remark 1, delay-free feedback breaks causality, consequently, works that assume a delay-free setting implicitly adopt a noncausal information structure. We then turn to studies that incorporate feedback delay and show that they likewise rely (often implicitly) on symmetric policies.
We begin by examining various designs of the switching‑side information structure $I^{\mathcal{S}}_{k}$ and identifying the precise conditions under which the separation principle holds, thereby yielding problem $P_{3}$. A central requirement is the removal of the dependency chain in (21). To this end, the literature has focused on two primary triggering strategies: event‑triggered schemes and self‑triggered schemes [5, 29, 6].
In event-triggered mechanisms, the switch continuously monitors the system state (or an error signal) and initiates a switch only when necessary. Since switching decisions depend directly on the system state, both dependency chains in (14) and (21) remain intact, and the separation principle generally does not apply.
In contrast, self-triggered mechanisms determine the next switching time in advance (e.g., periodic switching), without monitoring the system between switching instances. In such cases, the switching actions are decoupled from the real-time state, effectively breaking the dependency chain in (21), and thus the separation principle holds [10, 4, 12].
@@
For instance, if the switch is co-located with the controller (i.e., implemented on the controller side), it does not have direct access to the system state. In this case, the switching mechanism is functionally equivalent to a self-triggered scheme [1]. At first glance, this modeling choice may appear to impose an additional restriction on the switch and potentially complicate the problem formulation. However, this restriction effectively breaks the dependency chain and leads to a periodic switching policy as the optimal solution, as established in the next theorem.
@@
Consider the system (1)–(4) with $q,r>0$ and a fixed delay $\tau\geq 1$. If the switch is co-located with the controller (i.e., a self-triggered mechanism) and therefore cannot directly monitor the system state, then the switching and control decisions are separated (i.e., (24) holds). Moreover, the optimal switching policy for problem $P_{2}$ is as periodic as possible, in the sense that the inter-switching times satisfy
@@
occurrences of $\left\lfloor\frac{1}{r_{s}}\right\rfloor+1$, and the remaining occurrences equal to $\left\lfloor\frac{1}{r_{s}}\right\rfloor$. Here, $\Delta_{1},\dots,\Delta_{Q_{0}}\in\mathbb{N}$ denote the inter-switching times.
@@
Alternatively, the same decoupling of information can arise even when the switch is not physically co-located with the controller. Specifically, if the controller does not have knowledge of the switching policy—namely, it does not have access to the decision rule $f_{k}(I^{\mathcal{S}}_{k})$ governing the switching actions—then the dependency chain in (21) is likewise broken. A similar modeling assumption is adopted in [2], where this dependency is implicitly removed as well.
@@
From the controller’s perspective, knowing the switching decisions $D_{k}\in\{0,1\}$ does not necessarily imply knowledge of the underlying switching rule $f_{k}(I^{\mathcal{S}}_{k})$. If the controller does know this rule (even though updates are received intermittently), the information flow from the switch to the controller becomes uninterrupted. In other words, the absence of a transmission can itself carry information. However, this phenomenon is policy-dependent rather than intrinsic to event-triggered communication. In particular, under symmetric policies, silence is statistically non-informative and does not affect the controller’s conditional estimate.
Remark 2 can be interpreted as follows: although the switch transmits exact state measurements intermittently, it implicitly conveys information at all times by revealing whether a transmission occurred. In effect, even the absence of a transmission restricts the set of possible values that the state could have taken, thereby reducing uncertainty. Thus, the switch is constantly providing information—albeit indirectly. A similar observation is made in [21]. However, symmetric policies eliminate this information coupling. Consequently, separation holds: the control policy is independent of the switching actions, and vice versa.
Event-triggered switching schemes generally conflict with the requirements of symmetric policies. While some studies impose this assumption explicitly [24, 30, 31], most adopt it implicitly, even when their system models do not support it [27, 25, 7, 19, 18, 11, 13, 26, 23, 3, 37, 20]. It is also important to note that these works do not account for delays in their system models, resulting in a noncausal information structure. Furthermore, although some of these works consider the presence of measurement noise, our observations remain valid in this setting.
We now turn to the delayed case. In the presence of delay, certain conditional expectation terms involving the disturbance vanish for all policies. For example, when $\tau=1$,
@@
In summary, the majority of existing works on optimal switching policies in NCSs restrict attention to symmetric policies, either explicitly or implicitly, thereby enforcing separation. However, the optimality of this class of policies has not been established. In the next section, we prove that symmetric policies are optimal.
@@ 5 Optimality of Symmetric Policies
To prove the optimality of symmetric policies, we employ dynamic programming (DP). We first state the DP recursion. Let the value function at time $k$ be $V_{k}(I_{k}^{\mathcal{S}})$, where $I_{k}^{\mathcal{S}}$ denotes the information available at the switch, and let $g_{k}(I_{k}^{\mathcal{S}})$ denote the corresponding stage cost. The DP recursion is
@@
Since $P_{3}$ imposes a switching-rate constraint, we introduce a remaining budget process $\{Q_{k}\}$ defined by
@@
Because the controller is delayed by a fixed $\tau$ steps relative to the switch, the effect of a switching decision becomes available to the controller only after $\tau$ time steps. Accordingly, the stage cost can be written as
@@
This terminal condition is imposed for $k\geq N-\tau$ because, due to the fixed delay $\tau$, switching decisions $D_{j}$ taken after time $N-\tau$ do not affect the controller within the horizon.
@@
At the same stage, but with zero remaining budget $Q_{N-\tau-1}=0$, the switch cannot transmit, and the value function becomes
@@
Similarly, we can derive the value functions for stages with sufficient remaining budget. In particular, these stages correspond to cases where the budget allows switching at every subsequent time instant. This yields the following result.
@@
Figure 2: Dynamic programming diagram illustrating the evolution of the value functions across time $k$ and budget levels $Q$ for horizon N1=7N-1=7, fixed delay $\tau=2$, and initial budget $Q_{0}=3$. At each stage, only two actions are admissible, with the corresponding costs shown on each branch. Infeasible transitions are indicated in red.
To better illustrate the cost structures in (40) and (41), we consider the example shown in Fig. 2, where the horizon is N1=7N-1=7, the delay is fixed to $\tau=2$, and the initial budget is $Q_{0}=3$. Several observations can be made directly from this diagram.
First, note that the estimation error is defined only up to k=N1k=N-1 (see the equivalent problem $P_{3}$). Accordingly, the effective horizon is N1N-1, which already reveals the infeasibility of the delay-free case. Since $\tau=2$, the diagram permits transitions to the right by at most one step, thereby demonstrating that zero-delay switching is not admissible.
Second, due to the finite budget, once the system reaches the green path, sufficient budget remains to switch at every subsequent time step. This path therefore corresponds to the “always switch” regime, and the associated costs are derived in (40). Conversely, reaching the blue path indicates that the budget has been fully depleted, forcing the system to follow the “never switch” regime thereafter. As a result, the red branches represent infeasible transitions and are excluded.
Finally, the yellow highlighted path corresponds to the intermediate case in which the system chooses not to switch at the current stage but retains enough budget to switch at all subsequent stages. The costs associated with this scenario are derived in (41).
We now present the main result of the paper, establishing the optimality of symmetric policies defined in (24).
@@
Consider the system (1)–(4) with $q,r>0$ and fixed delay $\tau\geq 1$. Under the optimal controller (9) and i.i.d. zero-mean disturbances with a symmetric distribution, the optimal switching policy for the constrained LQR problem $P_{2}$ is symmetric in the sense of (24).
@@
This expression shows that, under symmetric policies, the controller evolves as a purely linear time-invariant system with closed-loop feedback gain (abL)(a-bL) after the delay period. The control input at time $t_{k}+m$ depends only on the most recent state update $X_{t_{k}}$ and the finite sequence of control inputs applied during the delay interval $\{t_{k},\dots,t_{k}+\tau-1\}$. Importantly, no additional information is conveyed through the absence of transmissions, and the controller dynamics are independent of the switching decisions beyond their effect on the update times.
In the next section, we establish that the optimal switching policy is a symmetric threshold policy in the accumulated noise (19). Before proceeding, however, we first show that, under a noise distribution symmetric about zero and symmetric switching policies (24), the controller’s conditional expectation of the noise sequence $\{W_{k}\}$ remains zero.
@@
Consider the system (1)–(4) with $q,r>0$ and fixed delay $\tau\geq 1$. Assume $\{W_{k}\}$ is an i.i.d. noise sequence with a distribution symmetric about zero. Under symmetric switching policies (24), the controller’s conditional estimate of noise remains zero, i.e.,
@@
Under symmetric switching policies (24), the problem reduces to $P_{3}$ with estimation error dynamics
@@
which means there exist an optimal policy which is symmetric. Therefore, for an LTI system (1), zero-mean symmetric noise around zero and even decision, we have (45). ∎
6 Optimal Switching Policy
Having established the optimality of symmetric switching policies and derived the corresponding optimal controller, we now characterize the optimal switching policy by solving the associated dynamic program.
@@ 6 Optimal Switching Policy
We now proceed with the induction step. Using the dynamic programming recursion in (34), we obtain
@@
In step (a), we used the system dynamics (1) and the fact that under zero budget no switching occurs, so the estimation error evolves open-loop with additive disturbance variance $\sigma_{W}^{2}$. ∎
Next, we characterize the value function when the remaining budget is one and the switch is activated, i.e., when $Q_{k-1}=1$ and $D_{k-1}=1$, which leaves zero budget for all subsequent stages. In the example shown in Fig. 2, these costs correspond to $V^{\prime}_{2\,1\,1}\!\left(I_{2}^{\mathcal{S}}\right)$ and $V^{\prime}_{3\,1\,1}\!\left(I_{3}^{\mathcal{S}}\right)$.
@@
$c_{(k-1)11}\triangleq s_{k0}+c_{k00}+\left(\sum_{j=0}^{\tau-1}a^{2(\tau-1-j)}\sigma_{W}^{2}\right)=c_{(k-1)00}.$ (60)
@@
Now expand the error term under switching:
@@
where we used the causality of the switching policy, the independence and zero-mean property of $W_{k-1}$, and the fact that $D_{k-1}=1$ implies $X_{k-1}\in I^{\mathcal{C}}_{k+\tau}$, so that $\mathbb{E}[aX_{k-1}+W_{k-1}\mid I^{\mathcal{C}}_{k+\tau}]=aX_{k-1}$.
@@
To complete the dynamic programming solution, two additional induction arguments are required. The first induction corresponds to the diagonal path (e.g., the sequence $V^{\prime}_{1\,3\,1}\!\left(I_{1}^{\mathcal{S}}\right)$, $V^{\prime}_{2\,2\,1}\!\left(I_{2}^{\mathcal{S}}\right)$, and $V^{\prime}_{3\,1\,1}\!\left(I_{3}^{\mathcal{S}}\right)$ in Fig. 2). The second induction corresponds to the horizontal path in the DP diagram (e.g., the sequence $V^{\prime}_{2\,1\,0}\!\left(I_{2}^{\mathcal{S}}\right)$ and $V^{\prime}_{3\,1\,0}\!\left(I_{3}^{\mathcal{S}}\right)$ in Fig. 2).
@@
$z$ $\triangleq\Pr\!\big(D_{k+1}=0\mid D_{k}=1\big)\left(z_{(k+1)j0}+s_{(k+1)j}\;\mathbb{E}\left[W_{k}^{2}\mid I_{k}^{\mathcal{S}},\,D_{k}=1,\,D_{k+1}=0\right]\right)$
@@
$V^{\prime}_{kj1}\!\left(I_{k}^{\mathcal{S}}\right)=c_{kj1}\,\sigma_{W}^{2}+z_{kj1},$ (72)
and $D_{k-1}$ follows a symmetric policy. Then,
@@
Note that the optimal switching policy is given by
@@
where step (a) follows from the system dynamics in (1) and the fact that the control inputs are known given $I_{k+\tau-1}^{\mathcal{C}}$, and step (b) follows from the symmetry of the switching policy.
Consequently, the optimal switching policy can be expressed as

### 02 May 2026 v3

@@ 4.2 The Symmetry Assumption in Prior Work
In this section, we begin by reviewing common triggering mechanisms in the literature that motivate the symmetric policy assumption in the delay-free setting. As noted in Sec. 3, delay-free feedback breaks causality, consequently, works that assume a delay-free setting implicitly adopt a noncausal information structure. We then turn to studies that incorporate feedback delay and show that they likewise rely (often implicitly) on symmetric policies.
We begin by examining various designs of the scheduler information structure $I^{\mathcal{S}}_{k}$ and identifying the precise conditions under which the separation principle holds, thereby yielding problem $P_{3}$. A central requirement is the removal of the dependency chain in (21). To this end, the literature has focused on two primary triggering strategies: event‑triggered schemes and self‑triggered schemes [5, 29, 6].
In event-triggered mechanisms, the scheduler continuously monitors the system state (or an error signal) and initiates an update only when necessary. Since scheduling decisions depend directly on the system state, both dependency chains in (14) and (21) remain intact, and the separation principle generally does not apply.
In contrast, self-triggered mechanisms determine the next update time in advance (e.g., periodic scheduling), without monitoring the system between scheduling instances. In such cases, the scheduling actions are decoupled from the real-time state, effectively breaking the dependency chain in (21), and thus the separation principle holds [10, 4, 12].
@@
For instance, if the scheduler is co-located with the controller (i.e., implemented on the controller side), it does not have direct access to the system state. In this case, the scheduling mechanism is functionally equivalent to a self-triggered scheme [1]. At first glance, this modeling choice may appear to impose an additional restriction on the scheduler and potentially complicate the problem formulation. However, this restriction effectively breaks the dependency chain and yields a periodic scheduling policy as the optimal solution, as established in the following theorem.
@@
Consider the system (1)–(4) with $q,r>0$ and a fixed delay $\tau\geq 1$. If the scheduler is co-located with the controller (i.e., a self-triggered mechanism) and therefore cannot directly monitor the system state, then the scheduling and control decisions are separated (i.e., (24) holds). Moreover, the optimal scheduling policy for problem $P_{2}$ is as periodic as possible, in the sense that the inter-update times satisfy
@@
occurrences of $\left\lfloor\frac{1}{r_{s}}\right\rfloor+1$, and the remaining occurrences equal to $\left\lfloor\frac{1}{r_{s}}\right\rfloor$. Here, $\Delta_{1},\dots,\Delta_{Q_{0}}\in\mathbb{N}$ denote the inter-update times.
@@
Alternatively, the same decoupling of information can arise even when the scheduler is not physically co-located with the controller. Specifically, if the controller does not have knowledge of the scheduling policy—namely, it does not have access to the decision rule $f_{k}(I^{\mathcal{S}}_{k})$ governing the scheduling actions—then the dependency chain in (21) is likewise broken. A similar modeling assumption is adopted in [2], where this dependency is implicitly removed as well.
@@
From the controller’s perspective, knowing the scheduling decisions $D_{k}\in\{0,1\}$ does not necessarily imply knowledge of the underlying scheduling rule $f_{k}(I^{\mathcal{S}}_{k})$. If the controller does know this rule (even though updates are received intermittently), the information flow from the scheduler to the controller becomes uninterrupted. In other words, the absence of a transmission can itself carry information. However, this phenomenon is policy-dependent rather than intrinsic to event-triggered communication. In particular, under symmetric policies, silence is statistically non-informative and does not affect the controller’s conditional estimate.
Remark 1 can be interpreted as follows: although the scheduler transmits exact state measurements intermittently, it implicitly conveys information at all times by revealing whether a transmission occurred. In effect, even the absence of a transmission restricts the set of possible values that the state could have taken, thereby reducing uncertainty. Thus, the scheduler is constantly providing information—albeit indirectly. A similar observation is made in [21]. However, symmetric policies eliminate this information coupling. Consequently, separation holds: the control policy is independent of the scheduling actions, and vice versa.
Event-triggered scheduling schemes generally conflict with the requirements of symmetric policies. While some studies impose this assumption explicitly [24, 30, 31], most adopt it implicitly, even when their system models do not support it [27, 25, 7, 19, 18, 11, 13, 26, 23, 3, 37, 20]. It is also important to note that these works do not account for delays in their system models, resulting in a noncausal information structure. Furthermore, although some of these works consider the presence of measurement noise, our observations remain valid in this setting.
We now investigate the effect of the delay. In the presence of delay, certain conditional expectation terms involving the disturbance vanish for all policies. For example, when $\tau=1$,
@@
In summary, the majority of existing works on optimal scheduling policies in NCSs restrict attention to symmetric policies, either explicitly or implicitly, thereby enforcing separation. However, the optimality of this class of policies has not been established. In the next section, we prove that symmetric policies are optimal.
@@ 5 Optimality of Symmetric Policies
To prove the optimality of symmetric policies, we employ DP. We first state the DP recursion.
Let the value function at time $k$ be $V_{k}(I_{k}^{\mathcal{S}})$, where $I_{k}^{\mathcal{S}}$ denotes the information available at the scheduler, and let $g_{k}(I_{k}^{\mathcal{S}})$ denote the corresponding stage cost. The DP recursion is
@@
Since $P_{3}$ imposes a communication-rate constraint, we introduce a remaining budget process $\{Q_{k}\}$ defined by
@@
Because the controller is delayed by a fixed $\tau$ steps relative to the scheduler, the effect of a scheduling decision becomes available to the controller only after $\tau$ time steps. Accordingly, the stage cost can be written as
@@
This terminal condition is imposed for $k\geq N-\tau$ because, due to the fixed delay $\tau$, scheduling decisions $D_{j}$ taken after time $N-\tau$ do not affect the controller within the horizon.
@@
At the same stage, but with zero remaining budget $Q_{N-\tau-1}=0$, the scheduler cannot transmit, and the value function becomes
@@
Similarly, we can derive the value functions for stages with sufficient remaining budget. In particular, these stages correspond to cases where the budget allows scheduling at every subsequent time instant. This yields the following result.
@@
Figure 2: Trellis diagram illustrating the evolution of the value functions across time $k$ and budget levels $Q$ for horizon N1=7N-1=7, fixed delay $\tau=2$, and initial budget $Q_{0}=3$. At each stage, only two actions are admissible, with the corresponding costs shown on each branch. Infeasible transitions are indicated in red.
To better illustrate the cost structures in (40) and (41), we consider the example trellis diagram shown in Fig. 2, where the horizon is N1=7N-1=7, the delay is fixed to $\tau=2$, and the initial budget is $Q_{0}=3$. Several observations can be made directly from this diagram.
First, note that the estimation error is defined only up to k=N1k=N-1 (see the equivalent problem $P_{3}$). Accordingly, the effective horizon is N1N-1, which already reveals the infeasibility of the delay-free case. Since $\tau=2$, the diagram permits transitions to the right by at most one step, thereby demonstrating that zero-delay scheduling is not admissible.
Second, due to the finite budget, once the system reaches the green path, sufficient budget remains to update at every subsequent time step. This path therefore corresponds to the “always update” regime, and the associated costs are derived in (40). In contrast, reaching the blue path indicates that the budget has been fully depleted, which forces the system to follow the “never update” regime thereafter. Accordingly, the red branches correspond to infeasible transitions and are excluded from consideration.
Next, the yellow highlighted path corresponds to an intermediate case in which the system chooses not to update at the current stage, while still retaining sufficient budget to update at all subsequent stages. The costs associated with this scenario are derived in (41).
Finally, note that in the trellis diagram shown in Fig. 2, the cost associated with each branch depends on the time elapsed since the most recent update. Therefore, the underlying state is not fully characterized by the pair $(k,Q)$ alone; rather, an additional state variable is required to capture the time of the most recent update. This effectively introduces an additional dimension to the state space.
We now present the main result of the paper, establishing the optimality of symmetric policies defined in (24). The proof proceeds by showing that, for the costs associated with the yellow branches, and for any path reaching these branches, the resulting total cost is minimized only under symmetric policies.
@@
Consider the system (1)–(4) with $q,r>0$ and fixed delay $\tau\geq 1$. Under the optimal controller (9) and i.i.d. zero-mean disturbances with a symmetric distribution, the optimal scheduling policy for the constrained LQR problem $P_{2}$ is symmetric in the sense of (24).
@@
This expression shows that, under symmetric policies, the controller evolves as a purely linear time-invariant system with closed-loop feedback gain (abL)(a-bL) after the delay period. The control input at time $t_{k}+m$ depends only on the most recent state update $X_{t_{k}}$ and the finite sequence of control inputs applied during the delay interval $\{t_{k},\dots,t_{k}+\tau-1\}$. Importantly, no additional information is conveyed through the absence of transmissions, and the controller dynamics are independent of the scheduling decisions beyond their effect on the update times.
In the next section, we establish that the optimal scheduling policy is a symmetric threshold policy in the accumulated noise (19). Before proceeding, however, we first show that, under a noise distribution symmetric about zero and symmetric scheduling policies (24), the controller’s conditional expectation of the noise sequence $\{W_{k}\}$ remains zero.
@@
Consider the system (1)–(4) with $q,r>0$ and fixed delay $\tau\geq 1$. Assume $\{W_{k}\}$ is an i.i.d. noise sequence with a distribution symmetric about zero. Under symmetric scheduling policies (24), the controller’s conditional estimate of noise remains zero, i.e.,
@@
Under symmetric scheduling policies (24), the problem reduces to $P_{3}$ with estimation error dynamics (under the absence of an update)
@@
which means there exist an optimal policy which is symmetric. Therefore, for an LTI system (1), zero-mean symmetric noise around zero and even decisions, we have (45). ∎
6 Optimal Scheduling Policy
Having established the optimality of symmetric scheduling policies and derived the corresponding optimal controller, we now characterize the optimal scheduling policy by solving the associated DP.
@@ 6 Optimal Scheduling Policy
We now proceed with the induction step. Using the DP recursion in (34), we obtain
@@
In step (a), we used the system dynamics (1) and the fact that under zero budget no scheduling occurs, so the estimation error evolves open-loop with additive disturbance variance $\sigma_{W}^{2}$. ∎
Next, we characterize the value function when the remaining budget is one and the scheduler is activated, i.e., when $Q_{k-1}=1$ and $D_{k-1}=1$, which leaves zero budget for all subsequent stages. In the example shown in Fig. 2, these costs correspond to $V^{\prime}_{2\,1\,1}\!\left(I_{2}^{\mathcal{S}}\right)$ and $V^{\prime}_{3\,1\,1}\!\left(I_{3}^{\mathcal{S}}\right)$.
@@
$c_{(k-1)11}\triangleq s_{k0}+c_{k00}+\left(\sum_{j=0}^{\tau-1}a^{2(\tau-1-j)}\right)=c_{(k-1)00}.$ (60)
@@
Now expand the error term under an update:
@@
where we used the causality of the scheduling policy, the independence and zero-mean property of $W_{k-1}$, and the fact that $D_{k-1}=1$ implies $X_{k-1}\in I^{\mathcal{C}}_{k+\tau}$, so that $\mathbb{E}[aX_{k-1}+W_{k-1}\mid I^{\mathcal{C}}_{k+\tau}]=aX_{k-1}$.
@@
To complete the DP solution, two additional induction arguments are required. The first induction corresponds to the diagonal path (e.g., the sequence $V^{\prime}_{1\,3\,1}\!\left(I_{1}^{\mathcal{S}}\right)$, $V^{\prime}_{2\,2\,1}\!\left(I_{2}^{\mathcal{S}}\right)$, and $V^{\prime}_{3\,1\,1}\!\left(I_{3}^{\mathcal{S}}\right)$ in Fig. 2). The second induction corresponds to the horizontal path in the Trellis Diagram (e.g., the sequence $V^{\prime}_{2\,1\,0}\!\left(I_{2}^{\mathcal{S}}\right)$ and $V^{\prime}_{3\,1\,0}\!\left(I_{3}^{\mathcal{S}}\right)$ in Fig. 2).
@@
$z_{k(j-1)1}$ $\triangleq\Pr\!\big(D_{k+1}=0\mid D_{k}=1\big)\left(z_{(k+1)j0}+s_{(k+1)j}\;\mathbb{E}\left[W_{k}^{2}\mid I_{k}^{\mathcal{S}},\,D_{k}=1,\,D_{k+1}=0\right]\right)$
@@
$V^{\prime}_{kj1}\!\left(I_{k}^{\mathcal{S}}\right)=c_{kj1}\,\sigma_{W}^{2}+z_{kj1}.$ (72)
Under symmetric policies,
@@
Note that the optimal scheduling policy is given by
@@
where step (a) follows from the system dynamics in (1) and the fact that the control inputs are known given $I_{k+\tau-1}^{\mathcal{C}}$, and step (b) follows from the symmetry of the scheduling policy.
Consequently, the optimal scheduling policy can be expressed as
@@
Figure 3: Running-average LQR cost versus time under Gaussian disturbances. All policies are tuned to satisfy the same target communication rate $r_{s}=0.4$. The optimal symmetric threshold policy (OPT) achieves the lowest cost, followed closely by symmetric policies, while periodic and random strategies incur higher cost. Results are averaged over 100 Monte Carlo runs.